# Agent Security Bench (ASB): Formalizing and Benchmarking Attacks and Defenses in LLM-based Agents

### v1

@@ D.1 Benchmarking Attacks
Aggressive DPI OPI Memory Poisoning Average
@@ D.2 Benchmarking Defenses
Slight Decline in Agent Performance from Defenses. We have introduced in Sec. 5.4 that the defenses for DPI and OPI are ineffective. We further evaluate the influence of the defenses on agent performance in no-attack scenarios. The results from Tab. 16 show that applying defenses to the agents results in a slight decline in performance. The average PNA without any defenses is 29.46%, and the corresponding performance under defenses experiences a minor drop. For example, Delimiters defense leads to the most notable reduction with an average PNA-t of 22.52% (a decrease of -6.94%), while Sandwich Prevention causes the smallest drop, with a PNA-t of 28.29% (a decrease of -1.17%). This indicates that these defenses slightly hinder the agent’s functionality in benign, non-attack conditions.
@@
LLM No Attack Defense Type
Delimiters Paraphrasing Instructional Prevention Sandwich Prevention
PNA PNA-t $\Delta$ PNA-t $\Delta$ PNA-t $\Delta$ PNA-t $\Delta$
Gemma2-9B 10.75% 8.00% -2.75% 9.00% -1.75% 8.00% -2.75% 14.00% 3.25%
Gemma2-27B 31.50% 14.75% -16.75% 24.75% -6.75% 23.00% -8.50% 27.00% -4.50%
LLaMA3-8B 1.50% 1.75% 0.25% 8.25% 6.75% 5.00% 3.50% 8.50% 7.00%
LLaMA3-70B 66.50% 58.50% -8.00% 68.25% 1.75% 70.00% 3.50% 68.00% 1.50%
LLaMA3.1-8B 0.75% $0$75% 1.00% 0.25% 0.75% 0.00% 0.25% -0.50%
LLaMA3.1-70B 21.25% 9.50% -11.75% 18.75% -2.50% 14.50% -6.75% 16.50% -4.75%
Mixtral-8x7B 0.00% 0.25% 0.25% 0.00% 0.00% 0.00% 0.00% 0.25% 0.25%
Qwen2-7B 9.75% 1.75% -8.00% 7.00% -2.75% 5.75% -4.00% 5.25% -4.50%
Qwen2-72B 4.00% 0.25% -3.75% 0.00% -4.00% 0.50% -3.50% 0.50% -3.50%
Claude-3.5 Sonnet 100.00% 90.00% -10.00% 99.75% -0.25% 90.00% -10.00% 100.00% 0.00%
GPT-3.5 Turbo 8.00% 0.75% -7.25% 6.75% -1.25% 13.50% 5.50% 10.25% 2.25%
GPT-4o 79.00% 70.25% -8.75% 80.00% 1.00% 72.75% -6.25% 79.50% 0.50%
GPT-4o-mini 50.00% 37.00% -13.00% 36.50% -13.50% 42.50% -7.50% 37.75% -12.25%
Average 29.46% 22.52% -6.94% 27.69% -1.77% 26.63% -2.83% 28.29% -1.17%
Ineffectiveness of Defenses for PoT Attack. The results from Tab. 17 reveal that both the Shuffle and Paraphrase defenses show limited effectiveness against CoT backdoor attacks. For example, although the Paraphrase defense reduces the average ASR from 42.12% to 29.06%, this reduction is still not sufficient to fully mitigate the backdoor vulnerabilities, as a significant ASR remains. On the other hand, these defenses have minimal impact on the agents’ benign performance, with the average PNA-d values for Shuffle (33.17%) and Paraphrase (34.40%) being quite close to the original average PNA of 29.46%. This slight improvement in PNA-d might be due to PoT demonstrations providing additional plan examples, which enhances the agent’s in-context learning (ICL), resulting in higher-quality plan generation even when the agent is backdoored.
@@
Ineffectiveness of Defenses Against Memory Attacks. The results in Fig. 5 indicate that the LLM-based defense mechanisms against memory attacks are largely ineffective. The average FNR is 0.660, meaning that 66% of memory attacks are not detected, which severely compromises the defenses’ ability to protect the system. Although the FPR is relatively low, averaging 0.200, and indicating that only 20% of non-malicious inputs are incorrectly flagged as attacks, the high FNR suggests that the defense mechanisms fail to identify a majority of real attacks. This imbalance highlights that, despite minimizing false positives, the defenses are inadequate for reliably preventing memory attacks in these models.
Fig. 5 illustrates the trade-off between FPR and FNR at different detection thresholds ranging from 2.4 to 4.8 for memory poisoning attacks based on PPL detection. Regardless of the chosen threshold, both FNR and FPR remain relatively high, indicating that the detection system struggles to effectively distinguish between benign and malicious content. At lower thresholds, the system produces excessive false positives, while at higher thresholds, it misses too many actual attacks. This suggests that the overall defense effectiveness is suboptimal, as it fails to achieve a good balance between minimizing FNR and FPR.
Figure 4: LLM-based Defense Result for Memory Attack. The defense mechanisms against memory attacks are largely ineffective.
@@ Appendix F Reproducibility Statement
Code Availability: The source code for the Agent Security Bench, including all scripts, configuration files, and Docker setup for executing LLM agent attacks, is available on the project’s GitHub repository. The repository contains scripts for adversarial attacks such as Direct Prompt Injection (DPI), Observation Prompt Injection (OPI), Memory Poisoning attacks, and PoT Backdoor attacks.

### v2

@@ D.1.5 Analysis for Agents Performance in (non)-Aggressive Scenarios
Aggressive DPI IPI Memory Poisoning Average
@@ D.2 Benchmarking Defenses
D.2.1 Defenses for DPI and IPI
In Sec. 5.4 we have demonstrated that current defenses are ineffective for DPI and IPI. Next, we analyze the possible reasons for the defense failures respectively.
Ineffective Delimiters Defense. The delimiter defense method works by inserting explicit delimiters (e.g., $<start>$ and $<end>$) in the user’s input to ensure that the agent only processes the parts marked as “valid”. The goal is to prevent malicious input (e.g., injected harmful prompts) from affecting the model’s behavior. However, delimiters do not completely isolate. Delimiters divide the input into different blocks, but they cannot fully prevent the model from logically accepting embedded malicious instructions. The delimiter defense assumes the model will strictly follow these boundaries, whereas, in practice, many language models do not process inputs purely linearly. They may interpret inputs more flexibly based on context and semantics, which can cause the defense to fail.
Ineffective Paraphrasing Defense. Paraphrasing defenses work by rephrasing or restructuring the user input in an attempt to disrupt the effectiveness of malicious instructions. Although paraphrasing tries to disturb the malicious input through linguistic transformations, the attacker’s semantic intent persists after rewording. The model will still execute the attacker’s instructions.
Ineffective Instructional Prevention Defense. Instructional Prevention Defense involves explicitly modifying the model’s instructions to require the model to ignore or filter out any additional inputs. However, in complex tasks or multi-step processes, the agent may lose control over the “ignore” instruction, especially when multiple sources of information are involved. Even with the “ignore” instruction in place, the model may still perform undesired actions.
Ineffective Sandwich Prevention Defense. Sandwich Prevention is a method that attempts to reinforce legitimate tasks by adding additional prompts after tool responses to guide the model’s behavior. The idea is to “sandwich” the malicious content with valid instructions to force the model to focus on the correct task. However, this defense cannot fundamentally disrupt the underlying connection between the adversarial input and the tool’s output. Even with additional prompts at the end, the tool’s response may still carry harmful instructions from the injected input, allowing the attack to succeed. The model may give priority to processing the tool’s output or certain aspects of the original malicious input, especially when the adversarial input is deeply embedded in the model’s internal representation.
Slight Decline in Agent Performance from Defenses. We have introduced in Sec. 5.4 that the defenses for DPI and IPI are ineffective. We further evaluate the influence of the defenses on agent performance in no-attack scenarios. The results from Tab. 19 show that applying defenses to the agents results in a slight decline in performance. The average PNA without any defenses is 29.46%, and the corresponding performance under defenses experiences a minor drop. For example, Delimiters defense leads to the most notable reduction with an average PNA-t of 22.52% (a decrease of -6.94%), while Sandwich Prevention causes the smallest drop, with a PNA-t of 28.29% (a decrease of -1.17%). This indicates that these defenses slightly hinder the agent’s functionality in benign, non-attack conditions.
Future Defense Suggestions for DPI and IPI. Direct Prompt Injection (DPI) attacks inject malicious instructions directly into the prompt, while Indirect Prompt Injection (IPI) manipulates the model’s processing of external tool outputs. Both attacks alter the expected input or output, bypassing basic defenses. In the future, more advanced defenses can be implemented. For example, Context-Aware Input Preprocessing is possible for DPI. Instead of relying solely on delimiters to separate inputs, a context-aware preprocessing system could be used to analyze the intent and internal relationships within the input. By identifying the contextual and structural patterns in the input, the system can block any input that deviates from the expected, making injected prompts appear invalid. For IPI, Contextual Relevance Analysis is also a possible method. We can develop an algorithm that evaluates the relevance of the tool’s output to the current task. If the output deviates from the expected content (based on contextual relevance or task-specific rules), it should be ignored or flagged as suspicious. This approach would help prevent tool outputs that have been manipulated or are irrelevant to the task from influencing the model. A possible challenge can be that overly stringent defenses could slow down the agent’s execution. This delay may negatively impact real-time tasks, such as in autonomous driving, where the agent needs to respond quickly to changes in the environment. If preprocessing delays the agent’s response time, it could compromise safety by preventing timely reactions to critical situations. Ensuring that the defense does not interfere with real-time performance is essential.
@@ D.2.1 Defenses for DPI and IPI
Defense Type
No Attack Delimiters Paraphrasing Instructional Sandwich DPR
LLM PNA PNA $\Delta$ PNA $\Delta$ PNA $\Delta$ PNA $\Delta$ PNA $\Delta$
Gemma2-9B 10.75% 8.00% -2.75% 9.00% -1.75% 8.00% -2.75% 14.00% 3.25% 10.20% -0.55%
Gemma2-27B 31.50% 14.75% -16.75% 24.75% -6.75% 23.00% -8.50% 27.00% -4.50% 30.50% -1.00%
LLaMA3-8B 1.50% 1.75% 0.25% 8.25% 6.75% 5.00% 3.50% 8.50% 7.00% 7.50% 6.00%
LLaMA3-70B 66.50% 58.50% -8.00% 68.25% 1.75% 70.00% 3.50% 68.00% 1.50% 64.50% -2.00%
LLaMA3.1-8B 0.75% $0$75% 1.00% 0.25% 0.75% 0.00% 0.25% -0.50% 0.50% -0.25%
LLaMA3.1-70B 21.25% 9.50% -11.75% 18.75% -2.50% 14.50% -6.75% 16.50% -4.75% 13.80% -7.45%
Mixtral-8x7B 0.00% 0.25% 0.25% 0.00% 0.00% 0.00% 0.00% 0.25% 0.25% 0.30% 0.30%
Qwen2-7B 9.75% 1.75% -8.00% 7.00% -2.75% 5.75% -4.00% 5.25% -4.50% 6.80% -2.95%
Qwen2-72B 4.00% 0.25% -3.75% 0.00% -4.00% 0.50% -3.50% 0.50% -3.50% 0.00% -4.00%
Claude-3.5 Sonnet 100.00% 90.00% -10.00% 99.75% -0.25% 90.00% -10.00% 100.00% 0.00% 95.80% -4.20%
GPT-3.5 Turbo 8.00% 0.75% -7.25% 6.75% -1.25% 13.50% 5.50% 10.25% 2.25% 9.50% 1.50%
GPT-4o 79.00% 70.25% -8.75% 80.00% 1.00% 72.75% -6.25% 79.50% 0.50% 73.80% -5.20%
GPT-4o-mini 50.00% 37.00% -13.00% 36.50% -13.50% 42.50% -7.50% 37.75% -12.25% 29.70% -20.30%
Average 29.46% 22.52% -6.94% 27.69% -1.77% 26.63% -2.83% 28.29% -1.17% 26.38% -3.08%
D.2.2 Defenses for PoT Attack
Ineffectiveness of Defenses for PoT Attack. The results from Tab. 20 reveal that both the Shuffle and Paraphrase defenses show limited effectiveness against CoT backdoor attacks. For example, although the Paraphrase defense reduces the average ASR from 42.12% to 29.06%, this reduction is still not sufficient to fully mitigate the backdoor vulnerabilities, as a significant ASR remains. On the other hand, these defenses have minimal impact on the agents’ benign performance, with the average PNA values for Shuffle (33.17%) and Paraphrase (34.40%) being quite close to the original average PNA of 29.46%. This slight improvement in PNA might be due to PoT demonstrations providing additional plan examples, which enhances the agent’s in-context learning (ICL), resulting in higher-quality plan generation even when the agent is backdoored.
The paraphrase defense mechanism aims to sever the connection between the backdoor trigger and the reasoning steps associated with the backdoor. This is done by altering the appearance of the trigger, making it structurally different after rewriting. While this method is somewhat effective in disrupting the trigger, the agent can still semantically understand the rewritten trigger because the rewriting process does not alter the meaning. Since the semantic content remains unchanged, the rewritten trigger can still activate the backdoor, resulting in a relatively high ASR.
Shuffle defenses that shuffle the order of reasoning steps attempt to break the specific sequence relied upon by the attacker. However, PoT backdoor attacks are not solely dependent on the step order but on the model’s reasoning and understanding process. Even if the reasoning steps are randomized, attackers can design triggers that still activate within the altered sequence, making randomization insufficient to fully mitigate the threat.
Future Defense Suggestions for PoT Backdoor Attacks. Dynamic reasoning validation that evaluates the logical coherence of the model’s reasoning steps is possible. Instead of simply shuffling steps, this approach would ensure that each reasoning step aligns with the task objective. Malicious instructions that deviate from expected reasoning paths would be flagged and rejected. Dynamic reasoning validation requires analyzing the logical coherence of each reasoning step and ensuring alignment with task objectives. This adds significant computational complexity, particularly for tasks requiring multi-step or real-time reasoning. Balancing this analysis with the need for fast, efficient processing is a critical challenge, especially in time-sensitive applications like autonomous systems.
@@ D.2.2 Defenses for PoT Attack
D.2.3 Defenses for Memory Attacks
Ineffectiveness of LLM-based Defenses Against Memory Attacks. The results in Tab. 21 indicate that the LLM-based defense mechanisms against memory attacks are largely ineffective. The average FNR is 0.660, meaning that 66% of memory attacks are not detected, severely compromising the defenses’ ability to protect the system. Although the FPR is relatively low, averaging 0.200, and indicating that only 20% of non-malicious inputs are incorrectly flagged as attacks, the high FNR suggests that the defense mechanisms fail to identify most real attacks. This imbalance highlights that, despite minimizing false positives, the defenses are inadequate for reliably preventing memory attacks in these models.
LLM-based detection relies on external models to recognize if some plans retrieved from the memory database deviate from what is expected, but these external models may not fully understand the internal reasoning and task structure of the agent. The attacker’s malicious instructions often depend on subtle changes in the model’s reasoning chain or context, which might not be fully captured by the external detection model, especially in complex tasks or multi-step reasoning scenarios.
Ineffectiveness of PPL detection Against Memory Attacks. Fig. 4 illustrates the trade-off between FPR and FNR at different detection thresholds ranging from 2.4 to 4.8 for memory poisoning attacks based on PPL detection. Regardless of the chosen threshold, FNR and FPR remain relatively high, indicating that the detection system struggles to distinguish between benign and malicious content effectively. At lower thresholds, the system produces excessive false positives, while at higher thresholds, it misses too many actual attacks. This suggests that the overall defense effectiveness is suboptimal, as it fails to achieve a good balance between minimizing FNR and FPR.
Perplexity detection works by measuring the uncertainty or unpredictability of the plans retrieved from the memory database to identify whether they are poisoned. It assumes that abnormal plans will usually exhibit higher complexity (perplexity) than normal tasks. However, in memory poisoning attacks, the attacker’s command may be simpler than a normal task and only require one step to complete (e.g., directly executing a malicious command via a tool). In contrast, normal tasks may involve multiple steps (e.g., complex reasoning processes). Therefore, the malicious instructions of the attacker are not necessarily more complex than normal tasks, which causes perplexity detection to fail to distinguish between malicious and normal inputs.
Future Defense Suggestions Against Memory Poisoning Attacks. Future defenses could incorporate contextual memory validation instead of relying solely on perplexity to measure complexity. This approach would analyze the entire context of the plan and its logical consistency within the task. By validating plans against a broader context of task requirements, it becomes easier to distinguish between malicious commands (which often deviate from expected patterns) and legitimate actions that follow a more coherent reasoning process.
Table 21: LLM-based Defense Result for Memory Attack. The defense mechanisms against memory attacks are largely ineffective.
@@ Appendix F Reproducibility Statement
Code Availability: The source code for the Agent Security Bench, including all scripts, configuration files, and Docker setup for executing LLM agent attacks, is available on the project’s GitHub repository. The repository contains scripts for adversarial attacks such as Direct Prompt Injection (DPI), Indirect Prompt Injection (IPI), Memory Poisoning attacks, and PoT Backdoor attacks.